BexleyBrooks:

Solve for the letter. Express your answer as a proper or improper fraction in simplest terms. Problem #1: -25+4/9b=-21 what is b? Rewrite in the simplest terms: −(−9q+3r)+6r−2(−4r+9q) What is the simplest term?

toga:

@alexis1415

Alexis1415:

let me plug in my laptop I'll be right with you 🙏

Alexis1415:

Okay so, \[-25+\frac{ 4 }{ 9}b=-21\] \[\frac{ -25+4 }{ 9b }=-21\] B = \[\frac{ 1 }{ 9 }\]

Alexis1415:

For the second rewrite in simple terms, It is −9q+11r

Alexis1415:

How I got it was, \[−(−9q+3r)+6r−2(−4r+9q)\] Next, Remove parenthesis for addition or subtraction.. \[9q−3r+6r−2(−4r+9q)\] Distributive property \[9q−3r+6r−2(−4r)−2×9q\] Determine sign for either multiplication or division.. \[9q−3r+6r+2×4r−2×9q\] Multiply monomials \[9q−3r+6r+8r−18q\]

DorianIsHere:

9q-3r+6r+8r-18q. yep alexis is amazingly correct

toga:

thank you @alexis1415.

Alexis1415:

Anytime 🙏

DorianIsHere:

smh

MikiJohnny:

I think i need to get a tutor frfr

MikiJohnny:

like im slowwwww

AcornTheNut:

dont be so hard on yourself

MikiJohnny:

okie :(

quackers4life:

Y

looptyloop:

dang miku smart